# USCTNet: A deep unfolding nuclear-norm optimization solver for physically consistent HSI reconstruction

### v1

@@ C Low-Rank Subspace Proximal Mapping
(ii) Adaptive probing & threshold schedule. With $\Xi\!\sim\!\mathcal{N}(0,1)^{n\times s}$ and $G=\operatorname{Diag}(g)\,\Xi$, estimate the residual ratio:
@@
(iii) Subspace proximal & fusion. Project, pre-transform with a light convolution, and apply subspace SVT with learnable $\beta_{t+1}$:
$\widehat{U}_{k,t}=Q_{t}\,\mathrm{ConvSVT}_{\psi}\!\big{(}Q_{t}^{\top}U_{k};\beta_{t+1}\big{)}.$ (9)
Let $Q_{t}^{\top}U_{k}=\tilde{U}\Sigma V^{\top}$ be the thin SVD of the small matrix in $\mathbb{R}^{r\times n}$ and define
$\Sigma_{\beta}=\mathrm{softshrink}(\Sigma;\beta_{t+1}),\qquad\mathrm{ConvSVT}_{\psi}(\cdot)=\tilde{U}\,\Sigma_{\beta}\,V^{\top}.$ (10)

### v2

@@ C Low-Rank Subspace Proximal Mapping
(ii) Adaptive probing & threshold schedule. With $\Xi\!\sim\!\mathcal{N}(0,1)^{n\times s}$ and $G=\operatorname{Diag}(g)\,\Xi$, estimate the residual ratio, which is used as fusion weight in Eqn.10:
@@
(iii) Subspace proximal & fusion. At iteration $t$, $U_{k}$ is projected onto the subspace spanned by $Q_{t}$ (Eqn. 5) to form the compact matrix $B_{t}$. A depthwise-separable module, $\operatorname{ConvSVT}_{\psi}(\cdot)$, is then applied as a data-adaptive surrogate for singular-value shrinkage. Its output is combined with the identity via the gate $\alpha_{t}=\sigma(\beta_{t})\in(0,1)$, where $\beta_{t}$ denotes the threshold increment (Eqn. 8). The result is finally mapped back to the ambient space.
$\widetilde{B}_{t}=(1-\alpha)B_{t}+\alpha(\operatorname{ConvSVT}_{\psi}\!\big(B_{t})),\widehat{U}_{k,t}=Q_{t}\,{B}_{t}.$ (9)